# H line shape and radius-luminosity relation in 2.5D FRADO

### v1

@@ 4.2 Results
When focusing on the peak values of the time-delay histograms, we observed strong consistency with the observed H R-L relation. The slope of the radius-luminosity relation obtained from our peak time-delays aligns well with the empirical data, indicating that the FRADO model can accurately capture the distribution of material and the dynamic response of the BLR gas. Importantly, the spread in the predicted time-delays which is thought to be dependent to the accretion rate of the sources, also follows the trends observed in real AGNs, although the observed spread is much larger than what we see in our simulations. This suggests that peak time-delays can be the robust indicator of dynamics and geometry of BLR. In figure 8 we plot delays separately from near and far side of the distribution, separately for each viewing angle. The difference between the far and near side time delay is rising strongly with the viewing angle, which is expected. The near-side lags tend to fall below the standard R-L relation (Bentz et al., 2013) due to shorter light travel times, while far-side lags lie closer to or above it.
@@ 4.2.3 R-L relation
The observed R-L relation should combine all viewing angles but looking at the separate panels of figure 8 we see qualitatively that we will not produce the observed spread if we use all the models. Much shorter time delay for high Eddington ratio can be accommodated in the model when, selectively, we choose only near side of the BLR for high Eddington sources. It is not the same scenario as proposed by Wang et al. (2014a) where the disk is shielding the BLR region close to the equatorial plane and the outflow is practically conical, thus shortening the time delay both for close and far side. Here we would have to postulate that some medius covers the far side of the BLR. Such a mechanism is not built into FRADO.
@@ 5 Virial factor
The virial factor (commonly denoted as , also called correction factor) plays a critical role in estimating the mass of a supermassive black hole in AGNs using reverberation mapping techniques. Assuming a flattened disk-like BLR with predominantly Keplerian motion, the black hole mass, , is typically estimated using the virial equation
where is the radius of the broad line region (measured via reverberation mapping), is representative of velocity of particles in the BLR, is the gravitational constant, and is the virial factor which accounts for uncertainties related to the geometry, kinematics, and inclination of the BLR in AGNs. Modeling of the BLR in 2.5D FRADO lets us directly find the virial factor.
5.1 virial factor based on FWHM
If we consider the observed FWHM of the broad emission line to be representative of velocity range of particles in the BLR, one can find via
where is the observed FWHM which may be related to that of intrinsic through following (Collin et al., 2006), and is the inclination angle, the angle between the observer’s line of sight and the symmetry axis.
The results of dependence of , predicted by 2.5D FRADO on the black hole mass, Eddington ratio, viewing angle, and shape factor are presented in the figure 9. As obvious, the results imply a large departure of BLR from a virialized status. For low black hole masses low Eddington ratios, the starts at lower values and increases as these two parameters increase. There is a peak at black hole masses around for Eddington ratios around . For higher black hole masses, decreases again, particularly for higher accretion rates. The behavior of appears to depend strongly on both black hole mass and Eddington ratio. The results also show that decreases with increasing viewing angle for all of the three representative models, showing an anti-correlation, especially for the results based on the far-side time-delays. The shape factor also show that at higher inclination angles the line profiles tend to be more Gaussian-like while at lower angles Lorentzian-like profiles are favored.
@@
On the other hand, one can rather use the velocity dispersion of particles in BLR, , to compute the virial factor as
Figure 10 shows the dependence of on black hole mass. As can be seen, . The range of in bottom-right panel of the figure, though with a noticeable scatter for high black hole masses, spanning approximately from 0.5 to 2.5, is very well consistent with the general observational constraints of Onken et al. (2004); Collin et al. (2006) in which an squared sin factor is also included. The results in the top-left panel, if divided by 10, are also largely consistent with observational constraints, without introducing any inclination term. The other panels show that a strong segregation based on the viewing angle.
@@ 6 Discussion
The trends with the global parameters: black hole mass and Eddington ratio predicted by the FRADO model reproduce in an interesting way the trends observed in the data. Particularly the results for higher black hole masses (above ) are very promising. In case of very small black hole masses, the model seems to underestimate the intensity of the outflow. Below we address the successes and the shortcomings of the FRADO model in detail.
6.1 line profiles
@@
Higher mass black holes produce broader emission lines due to the higher orbital velocities of gas in the BLR. This is consistent with the virial theorem, which links the velocity dispersion of the BLR gas to the black hole mass (e.g., Peterson et al., 2004).
@@
While the accretion rate influences the overall intensity of the line emission, its effect on the profile shape in FRADO is less dramatic. Higher accretion rates tend to result in more asymmetric profiles, possibly due to winds or other outflow phenomena (e.g., Netzer, 2013; Elvis, 2000). As the accretion rate increases, radiation pressure can dominate and reduce the width of the line profile, often introducing asymmetry. At lower accretion rates, the disk is colder, the dust sublimation radius moves towards the black hole where the Keplerian rotation is faster, resulting in broader profiles due to higher orbital velocities in the BLR. This aligns with observational studies where AGNs with high accretion rates (near or above the Eddington limit) tend to have narrower profiles, while low accretion rate AGNs have broader lines due to less interference from such forces. FRADO also predicts the outflow, and its importance rises with the accretion rate, but this outflow is apparently underestimated for small black hole masses. This is seen in double-peak line profiles, predicted by FRADO for small black hole masses and high accretion rates, instead of observed single-peak Lorentzian shape for most Narrow Line Seyfert 1 galaxies. The line width is correct, which means that the launching radius is properly evaluated but the outflow or turbulence are not strong enough to remove the double-peak structure. We do not have a clear explanation of that. It may be that line driving plays a role when clouds are already launched but we need this extra force only for low mass objects. However, estimates of the role of the line driving force are beyond of the current paper. This would also require estimates of the filtering of the continuum through the inner High Ionization line region producing lines like HeII and CIV .
@@
The inclination of the BLR disk plays a crucial role in shaping the observed line profiles. Face-on views reveal the double-peaked structure expected from a rotating disk, while edge-on views produce broader and more asymmetric profiles. This is consistent with models of disk-like BLR geometries (e.g., Gaskell, 2009).
@@
As presented in the results, with the increase of accretion rate, the shape factor tends to decrease, especially for massive black holes, indicating a shift from more Gaussian profiles to Lorentzian-like profiles. This may be attributed to the increasing turbulence and velocity dispersion in the BLR gas at higher accretion rates (e.g., Sulentic et al., 2000). For higher black hole masses, the shape factor decreases, particularly at higher accretion rates. This suggests that the emission lines become broader and more Lorentzian-like for massive black holes, likely due to the stronger gravitational potential and more extreme velocity distributions in the BLR (e.g., Collin et al., 2006). The shape factor exhibits a minimum at intermediate viewing angles (), with Gaussian profiles being most prominent at low and high inclinations. This behavior can be linked to projection effects and the more complex geometry of the BLR with formation of an outflow structure at higher black hole masses, higher Eddington ratios, where the line of sight intersects regions of the BLR with higher velocity dispersion, leading to Lorentzian-like profiles. The shape factor is less sensitive to profile details than the presence or absence of a double-peak structure so the predicted trends are consistent with the discussion in Collin et al. (2006).
6.3 time-lags and R-L relation
@@
The slope predicted by FRADO for low Eddington ratio sources is slightly flatter than canonical slope 0.5 expected from analytical FRADO (Czerny & Hryniewicz, 2011) and the result by Bentz et al. (2013). However, this is not surprising, taking into account the newer results for the high Eddington ratio sources. The direct results from GRAVITY (GRAVITY Collaboration et al., 2024) which resolved BLR structure give a shallower slope than Bentz et al. (2013) and was consistent with higher accretion rates leading to shorter time-delays but the errors were large due to a small size of objects (four objects). Our model in its current version still underestimates the spread for high Eddington rate sources.
@@
There may be several mechanisms at work which could lead to shortening of the time delay. At higher accretion rates, the inner disk region may become denser and more opaque. This can preferentially obscure the far side of the BLR, particularly due to self-absorption or scattering (Netzer & Marziani, 2010; Wang et al., 2014b), or strong disk winds that can rise above the BLR and block the far side (e.g., Murray et al., 1995; Elvis, 2000). Also black hole spin affects the monochromatic flux and the ionization parameter in a different way which could also play a role (e.g. Czerny et al., 2019).
The near-side of BLR appears literally closer, more projected to the observer and thus more luminous (in flux per unit area), particularly at intermediate to high inclinations (e.g., Goad & Korista, 2014, 2015). Therefore the systematic scatter observed in the R–L relation for high-accretion-rate sources, can be attributed to the geometric and radiative weighting of the line response across the BLR. Specifically, our modeling suggests that this deviation correlates with a progressive dominance of near-side emission in the reverberation signal. But this geometrical effect can be applied to low accretors as well thereby we do see some of them also deviating from the standard R-L relation of Bentz et al. (2013), however at high viewing angles for which the closer side of the BLR is obscured for the observer by dusty torus, the far side of the BLR still can be visible for low-accretors due to lack of inner opaque disk winds and strong outflows while in the case of high accretors the far side is not visible at any angle. That’s why we can also find some observed low accretors located little above the standard R-L relation.
@@
Comparing our results in figure 8 with the observational data (Martínez-Aldama et al., 2019) indicates that the maximum viewing angle at which the H is observed is 39 degree, and more than this angle either the H flux becomes very dim or the transition to type 1.5 AGNs happens. The results also imply that the standard R-L relation calibrated for low-accretors (Bentz et al., 2013) mainly include those sources viewed at very low viewing angles. In the next paper, we will go though comprehensive detailed on implications of this model for the R-L relation and quasar cosmology.
@@
These results imply that accurate black hole mass estimations using the virial method based on FWHM need to account for many parameters including variations in the accretion rate, as it significantly influences the virial factor, especially in the intermediate and high-mass regimes.
@@
On the other hand, line dispersion, , is often considered a more reliable measure of the BLR kinematics than FWHM, as can be concluded from the figure 10. This is because it captures the full range of velocities within the BLR, including random motions and contributions from all components of the line profile. Studies such as Peterson et al. (2004) have emphasized that yields more accurate black hole mass estimates and reduces scatter in the calibration of the virial factor. FWHM tends to measure the velocity of the bulk of the gas, but it can simply miss the contributions from high-velocity wings, which may result in underestimation of the black hole mass. Additionally, FWHM is more sensitive to projection effects. As also implied from our results in 10, line dispersion may lead to smaller uncertainties in black hole mass estimates. For example, Collin et al. (2006) argued that using rather than FWHM reduces scatter in the relation too, the correlation between black hole mass and the velocity dispersion of the host galaxy’s bulge.
@@
There are important issues which are still not addressed in the current version of FRADO model. As we mentioned before, the vertical velocity for small black hole masses seem to be underestimated and we suspect it might be related to the line driving force. Our model launches clouds by the dust radiation pressure but perhaps line diving turns in as well, and we have no estimate of its role. Combing the two radiation pressure forces, however, is a challenge. The codes describing exclusively line driven winds exist (Risaliti & Elvis, 2010; Quera-Bofarull et al., 2023) but they have different architecture and combining them will not be a simple task. Incorporating the role of line driving may also affect our conclusion about the metallicity. We use here the assumption of the metallicity five times higher than solar since otherwise the launching force seems too weak.
@@
Finally, we do not specify what happens in the disk regions where the wind is not launched, and for low mass black holes this happens (in our computations) for low Eddington rates. In this case the disk surface can be directly irradiated by the central region which also could lead to formation of emission lines. On one hand, AGN disks are relatively flat and geometrically thin so direct irradiation may not be efficient (see e.g. Loska et al., 2004) but irradiation can be done by scattering in the highly ionized medium, for example warm absorber, as argued in the same work.
@@ 6.4 FRADO and microlensing
Since the 2.5D FRADO BLR model is able to reproduce a large variety of observed broad emission line profiles observed in AGN, it is also interesting to see if microlensing of the FRADO BLR can reproduce the line profile distortions observed in a number of gravitationally lensed AGN. A negative answer would imply a significant revision of the model.
Microlensing by stars in the lens galaxy of a lensed AGN is a phenomenon that can selectively magnify quasar subregions, in particular within the BLR, producing observable distortions in the emission line profiles that can be used to infer the size, geometry, and kinematics of the BLR. Based on generic BLR models, Hutsemékers & Sluse (2021), Hutsemékers et al. (2023, 2024a, 2024b), and Savić et al. (2024) have shown that Keplerian disk and equatorial wind models best describe the geometry and kinematics of the BLR, while polar wind models can be ruled out. Using the physically based modeling of the BLR in 2.5D FRADO with well-defined initial source parameters (thus avoiding arbitrary parameterization of the BLR structure), we should be able to simultaneously fit both the line profiles and their distortions to further investigate the BLR properties and hopefully directly infer source parameters such as the black hole mass and/or the accretion rate.
As a first step, we used the method described in Hutsemékers et al. (2023, 2024a, 2024b, and references therein), replacing the generic models with a 2.5D FRADO model of the BLR. We adopted a typical model with a black hole mass of 108 solar masses, an accretion rate of 0.1 in Eddington units and an inclination of 34°. The isovelocity maps generated by the 2.5D FRADO model were then convolved with microlensing magnification maps. For each position on the magnification map, a magnification profile was computed, this profile representing the wavelength-dependent distortion of the emission line profile induced by the microlensing effect ( indicates no distortion at all). The generated profiles were then selected to match a representative set of observed magnification profiles among those shown in (Hutsemékers et al., 2023, 2024a, 2024b; Savić et al., 2024). The observed profiles were scaled to a common velocity grid. Such profiles are shown in Fig. 11, demonstrating that microlensing of the 2.5D FRADO BLR model can reproduce the different line profile distortions observed in several lensed AGN; the blue and red wings can be either both magnified, both demagnified, or one magnified and the other demagnified. By comparing a large number of simulated line and magnification profiles with the observed ones, it would be possible to better constrain the BLR parameters than with generic models.

### v3

@@ 4.2.2 Peak time-delay option
When focusing on the peak values of the time-delay histograms, we observed a strong consistency with the observed H R–L relation. The slope of the R-L relation obtained from our peak time delays aligns well with the empirical data, indicating that the FRADO model can accurately capture the distribution of material and the dynamic response of the BLR gas. Importantly, the spread in the predicted time delays, which is thought to be dependent to the accretion rate of the sources, also follows the trends observed in real AGNs; although we note that the observed spread is much larger than what we see in our simulations. This suggests that peak time delays can be the robust indicator of BLR geometry.
In Figure 8, we plot the delays from the near and far sides of the distribution separately for each viewing angle. The difference between the far and near side time delays is increasing strongly with the viewing angle, which is expected. The near-side lags tend to fall below the standard R–L relation (Bentz et al., 2013) due to shorter light travel times, while far-side lags lie closer to or above it.
@@ 4.2.3 R-L relation
The observed R–L relation should represent an average over all viewing angles. However, by examining the separate panels in Figure 8, we can draw the qualitative conclusion that the observed spread cannot be reproduced if we include all models uniformly. The much shorter time delays observed for high Eddington ratio sources can be accommodated in the model only when we selectively consider the near side of the BLR for such objects.
We note that this is not the same scenario as the one proposed by Wang et al. (2014), where the disk is assumed to shield the BLR near the equatorial plane and the outflow adopts a conical geometry, thereby shortening the delay from both near and far sides. In our case, we have assumed that there is some medium, such as the inner outflow, obscuring the far side of the BLR.
@@ 5 Virial factor
The virial factor (commonly denoted as , also called correction factor) plays a critical role in estimating the mass of a supermassive black hole in AGNs using reverberation mapping techniques. Assuming a flattened disk-like BLR with predominantly Keplerian motion, the black hole mass, , is typically estimated using the virial equation,
where is the radius of the BLR (measured via reverberation mapping), is representative of velocity of particles in the BLR, and is the gravitational constant. The virial factor accounts for uncertainties related to the geometry, kinematics, and inclination of the BLR in AGNs. Modeling of the BLR in 2.5D FRADO enables us to directly determine the virial factor (as discussed below). However, we will devote a separate comprehensive study on constraining the virial factor from our model in the near future.
5.1 Virial factor based on the FWHM
If we consider the observed FWHM of the broad emission line to be representative of velocity range of particles in the BLR, we can find via
where is the observed FWHM and is the inclination angle (the angle between the observer’s line of sight and the symmetry axis). The observed FWHM may be related to that of intrinsic through (Collin et al., 2006).
Our results for the dependence of , as predicted by the 2.5D FRADO model, on the black hole mass, Eddington ratio, viewing angle, and shape factor are presented in Figure 9. Obviously, the results imply a significant departure of the BLR from a virialized state. For low black hole masses and low Eddington ratios, the starts at smaller values and increases with both parameters. There is a peak at black hole masses around and Eddington ratios around . For higher black hole masses, decreases again, particularly at higher accretion rates. This behavior suggests a strong dependence of on both black hole mass and Eddington ratio.
These results also indicate that decreases with increasing viewing angle across all three representative models, indicating a clear anti-correlation, especially for time delays measured from the far side of the BLR. The shape factor results similarly indicate that at higher inclinations, line profiles tend to be more Gaussian-like, while lower inclinations favor more Lorentzian-like shapes.
@@ 5.2 Virial factor based on
Alternatively, we can also use the velocity dispersion of particles in BLR, , to compute the virial factor as
Figure 10 shows the dependence of on the black hole mass. As we can see, the range for in bottom-right panel of the figure, spanning approximately from 0.5 to 2.5 (although with a noticeable scatter for high black hole masses) is very consistent with the general observational constraints of Onken et al. (2004); Collin et al. (2006). The inclination-independent results in the top-left panel, show the same trend as well; however, these values are ten times larger than the observational constraints. The other panels display a strong segregation, which is based on the viewing angle.
@@ 6 Discussion
Here, we address the advantages and shortcomings of the FRADO model in detail. We start with the line profiles, described below.
@@ 6.0.1 Black hole mass and line widths
Higher mass black holes produce broader emission lines due to the higher orbital velocities of gas in the BLR. This is consistent with the virial theorem, which links the velocity dispersion of the BLR gas to the black hole mass (e.g., Peterson et al., 2004). However, in high Eddington ratio sources, radiation pressure may also contribute significantly to the gas dynamics, potentially modifying the line widths and inferred masses (e.g., Marconi et al., 2008).
@@ 6.0.2 Accretion rate and line profile shape
While the accretion rate is expected to influence the total flux of the line emission, its effect on the profile shape in FRADO is also significant. Higher accretion rates tend to result in more asymmetric profiles, possibly due to winds or other outflow phenomena (e.g., Netzer, 2013; Elvis, 2000). As the accretion rate increases, radiation pressure can dominate and reduce the width of the line profile, often introducing asymmetry. At lower accretion rates, the disk is colder, and the dust sublimation radius moves closer to the black hole, where Keplerian rotation is faster. This results in broader profiles due to higher orbital velocities in the BLR. This outcome is aligned with observational studies, whereby AGNs with high accretion rates (near or above the Eddington limit) tend to have narrower profiles, while low-accretion-rate AGNs display broader lines due to less interference from such forces (Pounds et al., 1995; Du et al., 2016).
FRADO also predicts outflows and their importance rises with accretion rate, but the outflow appears to be underestimated for small black hole masses. This is evident in the double-peaked line profiles predicted by FRADO for low black hole masses and high accretion rates, in contrast to the observed single-peaked Lorentzian shapes in most Narrow Line Seyfert 1 galaxies.
The line width is consistent with observation, which suggests that the launching radius is properly evaluated; however, the outflow or turbulence is not strong enough to remove the double-peaked structure. We do not have a clear explanation for this. It may be that line driving plays a role once the clouds are already launched, but this extra force is only needed for low-mass objects in order to be consistent with observation. However, estimating the role of the line-driving force is beyond the scope of the current paper. This would also call for the filtering of the continuum through the inner HIL region producing lines such as HeII and CIV to be modeled.
@@ 6.0.3 Viewing angle and profile broadening
The inclination of the BLR disk plays a crucial role in shaping the observed line profiles. Face-on views reveal the double-peaked structure expected from a rotating disk, while edge-on views produce broader and more asymmetric profiles. This behavior is consistent with models of disk-like BLR geometries (e.g., Gaskell, 2009).
@@ 6.1 Shape factor
As presented in the results, with increasing accretion rate, the shape factor tends to decrease, especially for massive black holes, indicating a shift from more Gaussian-like to Lorentzian-like profiles. This may be attributed to an increase in turbulence and velocity dispersion in the BLR gas at higher accretion rates (e.g., Sulentic et al., 2000). For higher black hole masses, the shape factor decreases further, particularly at high accretion rates. This suggests that the emission lines become broader and more Lorentzian-like for massive black holes, likely due to the stronger gravitational potential and more extreme velocity distributions in the BLR (e.g., Collin et al., 2006). The shape factor also exhibits a minimum at intermediate viewing angles (), with Gaussian profiles being more prominent at both low and high inclinations. This behavior may be linked to projection effects and the complex geometry of the BLR, particularly with the formation of an outflow structure at higher black hole masses and Eddington ratios, where the line of sight intersects regions of the BLR with higher velocity dispersion, leading to Lorentzian-like profiles. The shape factor is less sensitive to detailed profile features than the presence or absence of a double-peak structure; therefore, the predicted trends remain consistent with the discussion in Collin et al. (2006).
6.2 Time lags and the R-L relation
@@
The slope predicted by FRADO for high Eddington ratio sources is slightly flatter than the canonical slope of 0.5 expected from analytical FRADO (Czerny & Hryniewicz, 2011) and the result by Bentz et al. (2013). However, this is not surprising, taking into account newer findings for high Eddington ratio sources. The direct results from GRAVITY (GRAVITY Collaboration et al., 2024), which resolved the BLR structure, show a shallower slope than Bentz et al. (2013). This is consistent with higher accretion rates leading to shorter time delays, although the uncertainties were large due to the small number of observed objects (four). Our model, in its current version, still underestimates the deviation of high-accretion sources relative to the standard relation of Bentz et al. (2013), meaning that for a given viewing angle and selected visibility (whether near side, far side, or both), the difference in time delays between high- and low-accretion sources is not comparable to the observed spread of 0.2 to 0.4 dex (Bentz et al., 2013; Du et al., 2016; Du & Wang, 2019b; Mart´ınez-Aldama et al., 2019).
@@
Although the model does not predict the spread directly, several mechanisms could contribute to the shortening of the time delay. At higher accretion rates, the inner disk region may become denser and more opaque. This could preferentially obscure the far side of the BLR, due to self-absorption or scattering (Netzer & Marziani, 2010; Wang et al., 2014), or due to strong disk winds that rise above the BLR and block the far side (e.g., Murray et al., 1995; Elvis, 2000). Additionally, black hole spin affects the monochromatic flux and the ionization parameter in distinct ways, which could also influence the observed time delays (e.g., Czerny et al., 2019).
The near side of the BLR contributes more strongly to the observed emission-line response, particularly at moderate to high inclinations, due to reduced obscuration and more favorable radiative conditions compared to the far side. Therefore, the systematic scatter observed in the R–L relation for high accretion-rate sources can be attributed to geometric asymmetries and radiative weighting of the line response across the BLR. Specifically, our modeling suggests that this deviation correlates with a progressive dominance of near-side emission in the reverberation signal. Although this effect may also apply to low accretors, as some are seen to deviate from the standard R–L relation of Bentz et al. (2013). At high inclinations, where the near side of the BLR is obscured by the dusty torus, the far side can remain visible in low-accretion systems due to the absence of inner disk winds. In contrast, in high-accretion systems, inner outflows also end up obscuring the far side. This might explain why some observed low accretors are found slightly above the standard R–L relation. Referring to the possible partial presence of matter-bounded clouds (as we have considered in this work), such conditions could indeed reduce the far-side emissivity, particularly in high-Eddington sources. Thus, this effects could contribute to the diminished response. However, this effect is expected to be minor; otherwise, we would expect the near-side emission to be similarly suppressed.
@@
Comparing our results in Figure 8 with the observational data (Mart´ınez-Aldama et al., 2019) indicates that the maximum viewing angle at which H is typically observed is around 39∘; beyond this angle, either the H flux becomes very weak or the transition to type 1.5 AGNs occurs. The results also imply that the standard R–L relation, calibrated primarily for low accretors (Bentz et al., 2013), mostly includes sources observed at relatively low inclinations. In a forthcoming paper, we will present a more comprehensive analysis of the implications of this model for the R–L relation and its application to quasar cosmology.
@@
These findings suggest that accurate black hole mass estimates using the virial method based on FWHM must account for multiple parameters, including the accretion rate, as it has a significant impact on the virial factor, particularly in the intermediate to high-mass regime.
@@
On the other hand, the line dispersion, , is often considered a more reliable indicator of BLR kinematics than FWHM, as also supported by Figure 10. This is because captures the full range of velocities within the BLR, including contributions from both random motions and the extended wings of the emission line profile. Studies such as that of Peterson et al. (2004) have emphasized that yields more accurate black hole mass estimates and reduces the scatter in virial factor calibrations. While the FWHM primarily reflects the velocity of the bulk of the gas, it can overlook high-velocity components, leading to potential underestimation of black hole masses. Moreover, the FWHM is more sensitive to inclination and projection effects. As shown in our results in Figure 10, using line dispersion may lead to reduced uncertainties in black hole mass determinations. For instance, Collin et al. (2006) argued that adopting instead of FWHM also decreases the scatter in the relation, which connects black hole mass to the velocity dispersion of the host galaxy’s bulge.
@@
There are important issues that are not addressed in the current version of the FRADO model. As mentioned earlier, the vertical velocity for small black hole masses appears to be underestimated and we suspect this may be related to the absence of the line-driving force. Our model launches clouds via dust radiation pressure, but it is possible that line driving also plays a role; although we currently have no estimate for its contribution. Combining the effects of both radiation pressure mechanisms poses a significant challenge. Existing codes for exclusively line-driven winds (Risaliti & Elvis, 2010; Quera-Bofarull et al., 2023) are based on different frameworks and integrating them with our dust-driven model is not straightforward. Including the line-driving force could also impact our conclusions on the metallicity. In the current model, we adopted a metallicity that is five times higher than solar, as otherwise the launching force would become too weak.
@@
Finally, we have not specified what occurs in the disk regions where the wind is not launched. For low-mass black holes, this situation arises (in our computations) at low Eddington ratios. In such cases, the disk surface may be directly irradiated by the central source, potentially giving rise to emission lines. On one hand, AGN disks are relatively flat and geometrically thin, so direct irradiation may not be efficient (see e.g., Loska et al., 2004). However, irradiation could still occur via scattering in a highly ionized medium (e.g., a warm absorber) as also discussed in the cited work.
@@ 6.3 FRADO and microlensing
Since the 2.5D FRADO BLR model is able to reproduce a wide variety of broad emission line profiles observed in AGNs, it is also of interest to test whether microlensing of the FRADO BLR can account for the line profile distortions seen in a number of gravitationally lensed AGN. A negative result would imply the need for a significant revision of the model.
Microlensing by stars in the lensing galaxy of a lensed AGN can selectively magnify subregions of the quasar, particularly within the BLR. This leads to observable distortions in the emission line profiles, which can be used to probe the size, geometry, and kinematics of the BLR. Based on generic BLR models, Hutsemékers & Sluse (2021), Hutsemékers et al. (2023, 2024a, 2024b), and Savić et al. (2024) have shown that Keplerian disk and equatorial wind models best describe the geometry and dynamics of the BLR; whereas polar wind models can be excluded. Using the physically motivated 2.5D FRADO model, with well-defined source parameters and avoiding arbitrary parameterization, we should be able to simultaneously reproduce both the intrinsic line profiles and their microlensing-induced distortions, thereby providing new constraints on the BLR structure and potentially direct estimates of physical parameters such as the black hole mass and accretion rate.
As a first step, we used the method described in Hutsemékers et al. (2023, 2024a, 2024b, and references therein), replacing the generic BLR models with a 2.5D FRADO BLR model. We adopted a representative model with a black hole mass of , an accretion rate of 0.1 in Eddington units, and an inclination of 34°. The isovelocity maps generated by FRADO were then convolved with microlensing magnification maps. For each position on the magnification map, a magnification profile, , was computed; this profile represents the wavelength-dependent distortion of the emission line profile caused by microlensing ( where corresponds to no distortion). The resulting profiles were then compared with a representative set of observed magnification profiles taken from Hutsemékers et al. (2023, 2024a, 2024b); Savić et al. (2024), after scaling to a common velocity grid.
As illustrated in Fig. 11, microlensing of the 2.5D FRADO BLR model can reproduce the diverse line profile distortions observed in lensed AGNs. Both the blue and red wings of the lines can be magnified, demagnified, or affected asymmetrically. This result shows that by comparing a sufficiently large set of simulated line and magnification profiles to observed data, it could be possible to constrain BLR properties more effectively than with generic BLR models.